f(x) = x^2 – 4 and g(x) = x – 2
(a) What is (f + g)(x)?
(b) What is (f – g)(x)?
(c) What is (fg)(x)?
(d) What is (f/g)(x)?

Answer by Cromlix(4381)

Hi there,
a) (f + g)(x) = (x^2 - 4) + (x - 2)
= x^2 + x - 6
b) (f - g)(x) = (x^2 - 4) - (x - 2)
= x^2 - x - 2
c) (fg)(x) = (x - 2)^2 - 4
= x^2 - 4x + 4 - 4
= x^2 - 4x
d) (f/g)(x) = (x^2 - 4)/(x - 2)
= (x - 2)(x + 2)/(x - 2)
(x - 2)s top and bottom cancel out
leaving x + 2.
Hope this helps :-)